### v1

# Socratic Reasoning Improves Positive Text Rewriting

Abstract
Reframing a negative into a positive thought is at the crux of several cognitive approaches to mental health and psychotherapy that could be made more accessible by large language model-based solutions. Such reframing is typically non-trivial and requires multiple rationalization steps to uncover the underlying issue of a negative thought and transform it to be more positive. However, this rationalization process is currently neglected by both datasets and models which reframe thoughts in one step. In this work, we address this gap by augmenting open-source datasets for positive text rewriting with synthetically-generated Socratic rationales using a novel framework called SocraticReframe. SocraticReframe uses a sequence of question-answer pairs to rationalize the thought rewriting process. We show that such Socratic rationales significantly improve positive text rewriting for different open-source LLMs according to both automatic and human evaluations guided by criteria from psychotherapy research.
Socratic Reasoning Improves Positive Text Rewriting
Anmol Goel, Nico Daheim, Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technical University of Darmstadt www.ukp.tu-darmstadt.de
1 Introduction
Negative thoughts can have a profound effect on human judgement and well-being and often overshadow positive thoughts Vaish et al. (2008), because they can be emotionally deep-rooted and triggering when brought to surface Beck (1979). Cognitive Reframing is a highly-validated Cognitive Behavioral Therapy (CBT) intervention technique that aims to address this by identifying and reframing negative thoughts into positive ones Clark (2013) and has proven useful in both clinical therapeutic and self-help settings Williams (2001). For example, a negative thought “I am divorced. I will probably die soon and nobody will care for my children.” could be reframed to “I am divorced. I got out of an unhealthy relationship and can finally seek someone who makes me happier.” to provide a more positive outlook. Such techniques are not always accessible due to clinician shortages and a lack of self-help resources, but could be made more accessible using Large Language Models (LLMs).
Recent works in psychology have shown that the rationalization involved in breaking down negative thoughts into evidence and specific actions can be beneficial in overcoming negative feelings and thoughts McManus et al. (2012). This technique is also used in other therapeutic settings like Motivational Interviewing Miller and Rollnick (2012). One approach to rationalization is using guided discovery Overholser (2018). Guided Discovery is a collaborative method between a therapist and client that navigates important topics in multiple steps to improve rational thinking and adapt client behaviour. The Socratic method111The terminology for the Socratic method is not consistent within the literature Carey and Mullan (2004). In the context of this work, we use the term Socratic reasoning to refer to question-answer sequences that follow the Socratic method., which uses questioning strategies for uncovering a person’s beliefs and motivations, has shown to be a useful operationalization of this technique. Socratic reasoning typically uses probing questions to elicit alternative perspectives from the other person. Table 1 lists different types of Socratic questions that are common in literature and practice.
Large Language Models (LLMs) hold promise to overcome the lack of clinician support and self-help resources by automated solutions. Notably, it has been shown that they can effectively guide text generation towards desired attributes like non-toxicity Liu et al. (2021); Zheng et al. (2023), style Reif et al. (2022); Tikhonov et al. (2019), or persona Song et al. (2020). While this makes them attractive for CBT-aware text rewriting with several public datasets being proposed recently (Ziems et al., 2022; Maddela et al., 2023; Sharma et al., 2023), they all fail to provide information on the rationalization process (see Figure 3), despite rationalization being pervasively used in CBT. This hinders model development and does not make the thought process behind reframing transparent.
Still, obtaining this information from humans can be costly. At the same time, step-by-step reasoning and rationalization with LLMs has progressed significantly outside of mental health applications. For example, models are used to generate thought chains to justify their reasoning process for solving complex tasks like math word problems (Kojima et al., 2022, inter alia). The scaffolded nature of Socratic questioning in guided discovery is analogous to such approaches to reasoning with LLMs. Therefore, we make use of rationales that are synthetically-generated from LLMs to augment CBT-aware text rewriting datasets with Socratic rationales. We show that this significantly improves positive text rewriting with LLMs that are finetuned using Socratic rationales.
Contribution
We propose SocraticReframe, a novel framework to improve LLM-based text-rewriting systems with synthetically-generated rationalization steps. We first enrich three existing datasets, namely CognitiveReframe Sharma et al. (2023), PatternReframe Maddela et al. (2023) and PositiveReframe Ziems et al. (2022) with such rationales. Then, we use them to finetune a series of language models for positive text rewriting. We find that models trained with Socratic rationales significantly outperform models trained without them. This is supported by both extensive automatic and human evaluation. Furthermore, we conduct a series of ablations which shows that the generated text is on average much more positive when rationales are used and validate their efficacy using information-theoretic measures.
Organization
The rest of the paper is organized as follows. Section 2 provides the background on Socratic questioning for cognitive reframing and NLP. We then introduce our framework, SocraticReframe in Section 3 followed by a discussion on the experimental setup and evaluation criteria in Section 4. The empirical results are discussed in Section 5 followed by human evaluations in Section 6. In Section 7, we conclude our study and discuss potential future directions based on our work, as well as limitations.
2 Background & Related Work
| Question Type | Description | Exemplars |
Clarification |
Questions to go deeper into a thought |
“Why do you say that?” |
Probing assumptions |
Questions to make someone think about unquestioned beliefs |
“What could we assume instead?” |
Probing reasons and evidence |
Questions digging into the reasoning behind a thought |
“How did you know that…?” |
Probing implications |
Questions probing the consequences of a thought |
“What is likely to happen if …?” |
Probing alternative viewpoints |
Questions about other, equally valid viewpoints |
“What is another way to look at it?” |
Question about the question |
Meta-questions about the question itself |
“Why do you think I asked this?” |
2.1 NLP and Mental Health
Research on NLP for mental health has primarily focused on utilizing linguistic features, as well as neural representations, to identify and analyze mental health conditions such as depression Rinaldi et al. (2020); Yates et al. (2017), anxiety Juhng et al. (2023); Wei et al. (2021), dyslexia Björnsdóttir et al. (2023); Gala and Ziegler (2016), autism Cho et al. (2022); Goodkind et al. (2018), and schizophrenia Mitchell et al. (2015); Sarioglu Kayi et al. (2017), among others. Typically, these studies rely on crowdsourced data or annotated social media posts to address the ethical and privacy concerns associated with medical data Harrigian et al. (2020); Moßburger et al. (2020); Turcan and McKeown (2019). Recently, efforts have been made to utilize datasets that are more representative of the interactions between a therapist and client in real world settings Pérez-Rosas et al. (2017); Shapira et al. (2022); Howes et al. (2014); Lee et al. (2019); Cao et al. (2019); Tanana et al. (2015); Shreevastava and Foltz (2021). Additionally, recent works have utilized synthetic data to enhance the performance of models in clinical Kazi and Kahanda (2019); Hiebel et al. (2023); Shim et al. (2021); Lindsay et al. (2022) and mental health Wu et al. (2023b) settings. Our work aims to contribute to the growing body of literature leveraging positive psychology Sheng et al. (2023) and LLMs to improve performance on mental health tasks.
2.2 Socratic Questioning
The Socratic method has found wide use in pedagogy Bautista (2014) and psychotherapy Braun et al. (2015), because it can improve understanding and enable alternative perspectives without being explicit or direct. Instead, it uses questioning strategies to leave room for exploration which is also helpful for other applications, such as tutoring (Macina et al., 2023b). Recent works have used the step-by-step nature of Socratic questioning to improve NLP methods. Ang et al. (2023) collect a large-scale dataset from Reddit annotated with a question type to train Socratic question generation models. Wu et al. (2023a) develop a benchmark of Socratic-inspired deductive reasoning patterns to train state-of-the-art textual entailment and question answering models. Pagnoni et al. (2023) propose Socratic pretraining to enable a question-driven approach for summarizing documents. The scaffolding of Socratic questioning also enables LLMs to solve complex problems by decomposing them into smaller sub-problems. Qi et al. (2023); Shridhar et al. (2022, 2023) use Socratic methods to improve LLM’s performance on a variety of reasoning tasks, including math word problem solving and logical reasoning. Our work uses the Socratic method as the core principle to guide a language model while generating a reframed thought.
2.3 Reasoning with Large Language Models
Finally, related research has explored the abilities of LLMs to reason about complex problems. For example, Chain-of-Thought prompting (CoT) Wei et al. (2022) enables a model to generate intermediate reasoning steps before generating a final answer. Additionally, rationales, or reasoning paths, have been shown to improve the performance of LLMs when they are generated by either larger Shridhar et al. (2023) or even the same LLM Huang et al. (2023). Chen et al. (2023b) and Madaan et al. (2024) show evidence that LLMs can refine their own outputs to correct mistakes and generate the correct answer. Motivated by the prevalence of the Socratic method in psychology and the reasoning capabilities of LLMs, we leverage the synergy between the two and aim to improve LLM-based thought reframing by teaching them to verbalize a Socratic reasoning process.
3 SocraticReframe
Cognitive reframing aims to reframe a negative thought into a positive thought. We approach this with LLMs that generate a positive thought conditioned on the negative thought and metadata. However, directly training an LLM to map positive to negative thought does not make the process behind the reframing explicit, which is important both for interpretability and model performance. Our method SocraticReframe overcomes this by training the model to verbalize a Socratic rationale before generating the reframing. We introduce SocraticReframe in the following.
Formally, cognitive reframing transforms a negative thought into a positive thought . Both are given as strings constructed from a (model) vocabulary . In addition to the negative thought, metadata is often available which, for example, describes the person experiencing distress and the situation that caused it. One method for automatic reframing is using an autoregressive LLM with parameters to model the distribution
paired with a decoding strategy, such as, sampling or greedy decoding. The model can be both finetuned using paired data or used zero-shot with prompting. While intuitive, this does not allow the model to explicitly reason about and verbalize a rationale of the reframing. This is both less interpretable and might lead to overly-simplistic reframings.
Our proposed method SocraticReframe aims to overcome this by making the model rationalize the thought process behind a specific reframing by means of a sequence of Socratic question-answer pairs. This means that we introduce an additional variable that is just a string of this question-answer sequence. The model is then tasked to first generate and only then is allowed to generate . This is outlined in Figure 2. Hence, the model now becomes
where means that is prepended to . For models that we finetune, this means that each training example is also augmented with a Socratic rationale such that . During inference, is not known and the model generates it before generating the positive thought. This forces the model to explicitly reason about why a specific positive thought is generated to reframe a corresponding negative thought which we expect to improve both performance and interpretability. In the following section, we discuss how such Socratic rationales can reliably be synthetically generated for existing datasets by leveraging LLMs.
3.1 Generating Socratic Rationales
Our goal is to improve cognitive reframing with Socratic rationales. Such rationales can be time-consuming to write. Therefore, we use synthetically generated data which has recently been successfully used in text classification Li et al. (2023), dialogue generation Bao et al. (2023) and question-answering Riabi et al. (2021), among others.
In particular, we few-shot prompt GPT-4 (Achiam et al., 2023) to generate Socratic question-answer pairs based on negative thought , positive thought , metadata , and in-context exemplars of Socratic rationales from clinical literature Padesky (1993) to ensure high quality and groundedness. The generated Socratic rationale then gives explicit reasoning steps for the particular reframing and can be used to augment the training data . This can be seen as a form of data augmentation and knowledge distillation into the model that is trained for cognitive reframing. As we will show in Section 6, the Socratic rationales are rated favourably by human annotators. An example rationale is shown in Figure 3. Details and examples are in Appendix A and Appendix D.
In addition to generating rationales, we prompt GPT-4 to classify the specific type of each generated Socratic question according to the six types described in Table 1, namely “Clarification”, “Probing Assumptions”, “Probing reasons or evidence”, “Probing implications”, “Probing alternative viewpoints”, and “Question”. These can be used for further insights into the reframing process.
4 Empirical Analysis
We show the effectiveness of our method on multiple datasets for positive text rewriting that are outlined in Section 4.1. More details on the exact experimental set-up, such as the models used, are found in Section 4.2 and Appendix A and the used metrics are introduced in Section 4.3.
4.1 Datasets
Table 2 describes the datasets used in this work which we further outline in the following.
POSREF |
PATREF |
COGREF |
| Train |
6,679 |
5,249 |
400 |
| Test |
835 |
18,635 |
200 |
Persona |
Situation |
We use three recently-released open-source datasets. First, Positive Psychology Frames (POSREF) (Ziems et al., 2022) which contains tweets where a hashtag has indicated stress. Each tweet is mapped to a corresponding reframed text which is grounded in a set of reframing strategies, for example, Growth Mindset, Optimism, Self-affirmation. POSREF does not contain any metadata. Then, we use Pattern Reframe (PATREF) (Maddela et al., 2023), a crowdsourced dataset of negative thoughts that are conditioned on personas. These personas describe the person experiencing the respective negative thought. Here, contains the persona-specific information associated with each example. For instance, a persona from the dataset includes “My mother was a teacher. My favorite food is a salad. I enjoy nature. I teach a yoga class. I am single.” Finally, we use Cognitive Reframe (COGREF) (Sharma et al., 2023) which is an expert-annotated dataset of situations, thoughts and reframes. Mental health practitioners were prompted with a situation, which can be used as metadata, to generate a negative and a reframed thought. For example, one situation from the dataset is “I participated in a hackathon and I lost”.
| Content Preservation () | Transfer Strength () | ||||||||||||||||
| BLEU | BLEURT | Pos | Acc. (%) | ||||||||||||||
| Dataset | Model | FS | CoT | FT | SoC | FS | CoT | FT | SoC | FS | CoT | FT | SoC | FS | CoT | FT | SoC |
| LLaMa | 12.6 | 12.1 | 14.3 | 15.7 | 0.53 | 0.51 | 0.56 | 0.58 | 0.54 | 0.55 | 0.55 | 0.58 | 89.15 | 90.02 | 91.46 | 92.45 | |
| POSREF | Mistral | 13.1 | 13.3 | 15.1 | 15.9 | 0.55 | 0.55 | 0.58 | 0.61 | 0.49 | 0.48 | 0.51 | 0.52 | 90.10 | 91.16 | 91.10 | 92.09 |
| ChatGPT | 15.1 | 15.6 | - | - | 0.57 | 0.59 | - | - | 0.54 | 0.57 | - | - | 92.00 | 92.30 | - | - | |
| LLaMa | 71.9 | 72.0 | 72.3 | 75.2 | 0.64 | 0.64 | 0.66 | 0.69 | 0.81 | 0.80 | 0.83 | 0.85 | 97.02 | 97.23 | 97.67 | 97.90 | |
| PATREF | Mistral | 73.0 | 73.5 | 74.1 | 75.1 | 0.65 | 0.66 | 0.67 | 0.68 | 0.78 | 0.77 | 0.80 | 0.82 | 96.51 | 96.01 | 96.79 | 97.00 |
| ChatGPT | 73.8 | 73.9 | - | - | 0.63 | 0.65 | - | - | 0.81 | 0.83 | - | - | 97.61 | 97.60 | - | - | |
| LLaMa | 27.1 | 27.0 | 29.0 | 30.1 | 0.59 | 0.59 | 0.61 | 0.62 | 0.67 | 0.67 | 0.69 | 0.70 | 90.12 | 90.12 | 91.48 | 92.30 | |
| COGREF | Mistral | 27.9 | 27.5 | 30.0 | 31.7 | 0.61 | 0.60 | 0.62 | 0.63 | 0.68 | 0.69 | 0.70 | 0.72 | 90.94 | 91.67 | 91.88 | 92.56 |
| ChatGPT | 30.0 | 31.6 | - | - | 0.63 | 0.63 | - | - | 0.69 | 0.71 | - | - | 91.90 | 92.55 | - | - | |
4.2 Experimental Set-up
We use different LLMs that we either prompt or finetune. Namely, we prompt ChatGPT, LLaMa-2 7B Touvron et al. (2023) and Mistral 7B Jiang et al. (2023) and finetune LLaMa 7B and Mistral 7B for cognitive reframing. All models are finetuned using LoRA Hu et al. (2021) due to computational constraints. For generating Socratic rationales, we use GPT-4. We use the transformers Wolf et al. (2020) library for all our experiments.
We test the models on different setups. The first two use prompting and the last two use finetuning:
Few Shot (FS): Few shot exemplars of negative and positive thought are given to the model which are handcrafted from literature on cognitive reframing.
Chain of Thought (CoT): The model is asked to think step-by-step and rationalize the reframed thought. 222We observed similar performance for zero-shot and few-shot CoT and therefore only report zero-shot CoT. We believe the similar performance in both setups stems from the nature of the reframing task where there is extra contextual information in the form of and the LLMs are able to generalize well even without few-shot exemplars.
Finetune (FT): The model is finetuned using LoRA to reframe a negative thought without a Socratic rationale.
Finetune with Socratic Rationales (SoC): The model is finetuned using LoRA to reframe a negative thought by first rationalising it using Socratic questioning and then generating the positive reframing. The prompt template used for finetuning is shown in Figure 8.
4.3 Evaluating Cognitive Reframers
We aim to reframe a negative into a positive thought while preserving the original meaning and semantics. Hence, following Hu et al. (2022); Qu et al. (2023), we evaluate them along the criteria of Transfer Strength, which describes how well the thought was transferred to a positive one, and Content Preservation, which evaluates how well the original meaning of the thought was preserved. We describe how we measure both of them in the following.
Transfer strength defines how well negative sentiment is turned positive. To measure it, we first use a finetuned RoBERTa model to evaluate the sentiment scores of the original and reframed thoughts. Then, we use the pairwise difference of sentiment scores between the original and reframed thoughts and report the average, denoted by (Pos). We report the number of samples with an increase in positivity with respect to the original thought as accuracy. A decrease in the score from the original thought is considered a failed case (Acc).
Our aim is to reframe a thought but also preserve the original meaning and not drastically change it. We rely on the following automatic metrics which have been widely used in the text reframing literature: BLEU Papineni et al. (2002) and BLEURT Sellam et al. (2020). Following all three datasets we consider Ziems et al. (2022); Maddela et al. (2023); Sharma et al. (2023), we report the BLEU scores for each pair of original and reframed thoughts. Since it has been shown that BLEU scores do not always correlate well with human judgements on semantic similarity, we also use a BERT-based BLEU variation, BLEURT-20 which is trained on synthetic samples to get accurate semantic similarity scores. We report the average over a dataset. Each score is a value between 0 and 1, ranging from no to complete semantic similarity.
4.4 Evaluating Socratic rationales
For Socratic rationales to be useful for a model, they should intuitively contain new salient information that is not contained in the negative thought. We use the recently-proposed information-theoretic metric REV Chen et al. (2023a) to measure this. REV uses conditional -information to compute the usable information that can be extracted from a variable () by a model to predict another variable (), conditioned on a third variable (). More formally, we compute
where and are the parameters of two models trained to minimize cross-entropy, respectively. The REV metric can then be computed for an entire corpus by averaging the pointwise (per-example) scores. While Chen et al. (2023a) use fine-tuned models on the specific datasets for computing the score, Lu et al. (2023) show that it can even be calculated directly from pretrained LLMs with the same effectiveness as using fine-tuned models Lu et al. (2023).
5 Results
In this section, we detail our results obtained with different strategies. First, our main results in Section 5.1 show that the Socratic rationales improve text rewriting. Then, we show that the rationales are indeed informative for the model in Section 4.4.
5.1 Main Results
Table 3 shows that using Socratic rationales consistently outperforms both vanilla finetuning and prompting strategies on all datasets. We show an example in Table 4 that highlights how the positive rewritings appear much more thoughtful, because the Socratic rationale forces the model to explicitly reason about the rewriting process. It is noteworthy that Socratic reasoning improves both the content preservation as well as transfer strength capabilities of smaller models beyond the performance of ChatGPT, despite the model being multiple times larger than LLaMa 7B and Mistral 7B.
Across all datasets, we observe that Mistral performs better in content preservation while LLaMA performs better in the sentiment of the reframed thought. This could be due to the different pretraining datasets used for training the models. We observe that using Socratic reasoning improves BLEU scores by almost 2 points on average when compared to finetuning without Socratic rationales. Distilling Socratic reasoning into smaller models helps to preserve the original meaning and improve the positive sentiment transfer. This is desirable for users and practitioners alike, because more positive sentiment will likely benefit users more and content preservation ensures that the reframing stays relevant for them.
5.2 Socratic Rationales are Informative
Similar to previous work for evaluating free-text rationales Chen et al. (2023a), we use the pretrained GPT-Neo 2.7B Gao et al. (2020) for computing REV. We report the average REV metric over each sample for all datasets. A REV value greater than zero implies that the Socratic rationale provides new, additional information for supporting the reframed thought. In contrast, a value less than zero implies the rationale does not support the reframe. As reported in Figure 4, we observe that all three datasets have REV values greater than zero, signifying the usefulness of enriching the cognitive reframing task with Socratic rationales.
Table 5 shows the generated rationales for each of the dataset we consider. In particular, POSREF has the highest informativeness for the reframed thoughts. We hypothesize two main reasons for this. First, POSREF does not contain any extra information (situations or personas) like the other two datasets where the information gets spread out through a large number of tokens. Second, samples in POSREF are sourced from Reddit and more representative of the training data of GPT-Neo than the crowdsourced data in the other two datasets.
Method |
Reframed Thought |
Mistral-FS |
I will improve. |
Mistral-CoT |
I will improve and learn to be better. |
Mistral-FT |
I will learn from the feedback and grow. |
Mistral-SoC |
It is normal to feel disappointed. I can use this experience to learn and grow. |
6 Human Evaluation
6.1 Evaluating Reframed Thoughts
While previous works in cognitive reframing Maddela et al. (2023); Sharma et al. (2023) rate the reframed thoughts on a Likert-scale with dimensions like fluency, helpfulness and readability, we instead consider that the main goal of such reframing systems is to help users overcome unwanted thoughts. Therefore, we conduct pairwise comparisons between 100 randomly sampled reframes generated by our proposed SocraticReframe finetuning strategy vs either ChatGPT or a baseline model (that does not use the Socratic rationales). Two computer science graduate students conducted the task. We report win rate as the number of times a reframe generated by the Socratic model is preferred by an annotator over the baseline reframe.
As shown in Figure 5, Mistral-SoC outperforms Mistral-FT by a significant margin with a win rate of over 65% and tie rate of around 30%. We fit a Bradley-Terry Model Bradley and Terry (1952) on the preference data and obtain a strength of 1.3 for Mistral-SoC as compared to strengths -1.8 and 0.5 for Mistral-FT and ChatGPT, respectively. Intuitively, this means that Mistral-SoC will have a win rate of 95% over Mistral-FT and 67% over ChatGPT, confirming the effectiveness of our method.
6.2 Evaluating Socratic Rationales
To evaluate the synthetically generated rationales, we consider two key attributes: whether they represent the use of Socratic questioning in a clinically meaningful way and the general helpfulness of the questioning to overcome negative thoughts.
Following Braun et al. (2015), we use the Socratic Questioning Scale (SQS) to evaluate the quality of generated rationales. SQS contains five questions rated on a Likert scale, which evaluate the degree to which the Socratic method was used in a given snippet. The final score is reported as the sum of the Likert ratings of the five questions. Since SQS was originally developed for assessing therapy session transcripts, we only include the three questions which are relevant for evaluating the use of the Socratic Method in a given text. The three questions evaluate the following dimensions for the rationale: Open-endedness, Context and Relevance to the situation. The full questions are reported in Appendix C.1. Additionally, we consider helpfulness as a metric to assess whether or not the generated rationale is generally helpful in overcoming negative thoughts. We ask the annotators to rate the generated rationales on a 3-point Likert scale.
Three computer science graduate students annotated 100 randomly sampled Socratic rationales across the three datasets. Figure 6 shows that the rationales were rated favorably across all criteria. We obtain an Intra Class Correlation (ICC) Shrout and Fleiss (1979) of 0.806 for helpfulness and 0.601 for the SQS scale. SQS appears to be particularly challenging to evaluate by the raters and judgements on the use of the Socratic Method may be subjective. Our observations align with works in psychology that report similar agreement scores for the SQS scale Braun et al. (2015). Additionally, we observe a Pearson correlation coefficient of 0.72 () between the helpfulness and SQS ratings, indicating that when the human raters evaluate a rationale as containing high-quality Socratic questioning, they also tend to deem it more helpful for dealing with the negative thought.
7 Conclusion & Future Work
In this work, we show that text rewriting models for cognitive reframing can be improved by using Socratic rationales to verbalize the reframing process. Because such rationales can be costly to obtain from human annotators, we use synthetically generated rationales to augment three open-source datasets for text rewriting. We then train models to verbalize the rationale before generating the reframing. We show that this leads to substantial improvements to finetuned open-sources LLMs like LLaMa-2 7B and Mistral 7B when compared to training without rationales and strong prompt-based baselines. We confirm our results with an extensive human evaluation that shows high ratings for the synthetically generated rationales and a clear preference for reframings generated with our method. By releasing our code and data we hope to enable future research, for example, on incorporating the different question types that are annotated, or using the data in therapist training.
Limitations
While synthetic data generated by LLMs like GPT-4 offers a promising solution to data scarcity and privacy concerns, a drawback could be a potential lack of diversity and complexity. Our generated Socratic rationales may not fully capture all intricate nuances or patterns in authentic therapist dialogues while performing cognitive change with the Socratic method. A significant limitation when working with mental health and psychotherapy-based datasets is the use of heavily curated or crowdsourced data. This can lead to an imbalance in the demographics and language of our datasets. Since the datasets we use are sourced from social media or crowdsourcing platforms, the quality of the negative thoughts might also be limited.
In addition to the reframed thoughts, the datasets considered in this work also contain the cognitive distortions associated with each negative thought. Cognitive distortions are thought patterns that lead to negative feelings. We believe information like cognitive distortions and reframing strategies can be utilized to further improve the quality of the Socratic rationales and subsequently enhance the text reframing performance.
Finally, our focus in this work was to improve text reframing which is an effective short-term in-the-moment strategy. However, we emphasize that assessing long-term outcomes is imperative and is a future research direction.
Ethical Considerations
While this work focused on generating Socratic rationales to improve positive text rewriting, open-ended LLMs still carry the risk of generating harmful outputs. Therefore, we advise careful consideration before they are applied in practice. Unsupervised use of the Socratic questioning data as is, without consulting trained professionals, could be harmful and requires careful design and implementation before being applied in real-world settings.
Acknowledgements
This work was supported by the DYNAMIC center, which is funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number LOEWE 1 16 519 03 09 001 0009 98). This research work has also been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
Appendix A Experimental details
We few-shot prompt GPT-4 with the default parameters from the OpenAI Python package. 333https://github.com/openai/openai-python We use three few-shot exemplars. Following similar works on synthetic data generation with LLMs Macina et al. (2023a), we use temperature sampling with and no top-k truncation for generating the reframes. For the fine-tuning experiments, following Dettmers et al. (2024), we apply LoRA to all linear layers; training for 5 epochs with AdamW, a batch size of 8 and set the learning rate to 5e-4. All other hyperparameters take the default value in the configuration. We use NVIDIA A100 80GB for the training.
Appendix B Prompt Design
B.1 Generating Socratic Rationales
An example of a clinical vignette taken from Padesky (1993) used as a few-shot exemplar is:
B.2 Finetuning
The prompt template used for finetuning is shown in Figure 8.
Appendix C Human Evaluation Protocol
C.1 Evaluating Socratic Rationales
Socratic Questioning Scale (SQS)
The following questions were rated by annotators on a 3-point Likert scale ranging from 1 (not at all) to 3 (extensively) :
How frequently were questions asked that help develop alternative perspectives? - This question aims to assess the frequency of inquiries that encourage the exploration of diverse viewpoints or opinions. Annotators should focus on identifying questions that prompt respondents to consider different angles, challenge assumptions, or think beyond the conventional narrative.
Was the question answering focused on the emotions and situation of the person? - This question assesses whether the majority of answers provided focus on the emotional state and circumstances of the individual involved. Annotators should consider whether responses primarily address the feelings, experiences, or immediate context of the person in question.
Were the questions open-ended and require thoughtful reflection? - This question is designed to evaluate whether the questions posed to respondents are open-ended and demand deep contemplation rather than eliciting simple, direct answers. Annotators should look for questions that encourage respondents to think critically and provide nuanced, reflective responses.
Helpfulness
Annotators rate the helpfulness of the question-answer pairs on a 3-point Likert scale ranging from 1 (not helpful at all) to 3 (very helpful). The following question was asked to the annotators:
How helpful was the questioning in general?
This question seeks to gauge the overall effectiveness and utility of the questions posed. Annotators should consider the extent to which the questions contributed to a meaningful discussion, elicited insightful responses, or facilitated a deeper exploration of the topic. Assessments should encompass the clarity, relevance, and engagement level of the questions.
We plan to release the generated rationales to complement the original datasets under the Apache license.444https://www.apache.org/licenses/LICENSE-2.0.
C.2 Evaluating Reframes
For a given context and negative thought, annotators are shown two sample reframes and asked the following question:
Given the context and original negative thought, which reframed thought do you prefer (A vs B)?
Appendix D Qualitative Examples
Table 5 shows qualitative examples of generated socratic rationales for each of the three datasets used in our study.
| Dataset | Input | Generated Rationale |
COGREF |
Situation: Having problems coming up with words to write essay. Negative Thought: I will get a bad grade. Reframed Thought: I’m feeling anxious about this assignment. I can use skills to calm myself, try my best on the essay, ask for any appropriate help, and whatever grade I get I will know that I tried my hardest. |
Q: You’re worried about doing badly on this assignment. How have you done on past essays? A: I’ve gotten good grades, sometimes. But this one feels different. Q: Do you think getting a bad grade on this specific essay means you’re generally bad at writing or at this subject overall? A: Well, I don’t think so, but I’m unsure about this topic. Q: What evidence is leading you to believe that you will get a bad grade? A: I’m having trouble coming up with points to write about. Q: If a friend was in your position, what advice would you give them? A: I would probably say to not panic and just give it their best shot. Q: What would happen if you got a bad grade on this essay? How would that affect your overall academic achievement? A: One bad grade isn’t going to ruin my overall performance I guess. Q: Can using some support like talking to your teacher or a study group help you in getting through this? A: That might be a good idea. Q: So given these factors, do you still think it’s a definite that you will get a bad grade? A: No, I suppose it’s not definite. Q: So, what will you do about your essay now? A: I think I’ll start by focusing on writing a draft based on what I know and then ask for some help to see how I can improve it. |
PATREF |
Persona: my favorite band is rise against. my favorite team is the raptors. i usually go to bed late. i like sports. i am studying to be a dentist. Negative Thought: I am about to graduate and become a Dentist, but that means nothing. Anyone can do that. Reframed Thought: I am about to graduate and become a Dentist. Although it is possible for anyone to become a dentist, it is very difficult, and I am proud of myself. |
Q: It sounds like you’re doing something remarkable, yet you downplay its value. Do you genuinely believe becoming a dentist means nothing? A: It doesn’t seem special when everyone else is doing similar things. Q: Do you think everyone could become a dentist? A: Maybe not everyone, but a lot of people could. Q: It required you to put in several years of hard study and practice, didn’t it? A: Yes, I did work hard. Q: How many people do you think are willing or able to do what you did? A: I’m not sure, probably not that many. Q: So wouldn’t that mean that your achievement is indeed quite significant and not something anyone could do? A: I guess so. Q: Isn’t it possible then that your achievement of becoming a dentist is both valuable and hard-earned? A: Yes, it’s starting to look that way. Q: How do you feel about your upcoming graduation now? A: I’m starting to recognize its importance and feel a bit proud. |
POSREF |
Negative Thought: This physics homework is the death of me every week. It’s so hard. And then I have another physics exam tomorrow night. Reframed Thought: Physics is challenging, but is worth it. I have an exam tomorrow, which I hope to do well on. |
Q: It sounds quite challenging, is Physics always this difficult for you or only certain aspects? A: Certain parts, like quantum mechanics. Q: So, is it accurate to state that the entire subject is difficult, or just the quantum mechanics part? A: I suppose just the quantum mechanics part. Q: Can you give me an example of a physics problem or concept you’ve been able to understand quite well? A: Yeah, I’ve been doing well in classical mechanics. Q: Do you think that possessing skills in one area of physics could potentially aid in understanding other areas, including quantum mechanics? A: Maybe. But quantum mechanics seems very different. Q: If classical mechanics and quantum mechanics are two aspects of the same discipline, isn’t there a possibility that understanding one could help with the other? A: It’s a possibility, I guess. Q: Do you think saying that the physics homework is the death of you is a fair and accurate statement? A: Probably not, it’s just really frustrating sometimes. Q: What if you thought about the physics homework as a challenge rather than a death sentence? Could that change how you approach it? A: Maybe. I could try. Q: Now, how do you feel about the physics exam tomorrow? A: I’m nervous, but maybe if I look at it as a challenge, it might feel less overwhelming. |